# Unveiling the Basin-Like Loss Landscape in Large Language Models

### v2

@@ 6 Conclusion
In this work, we explore the loss landscape of large language models to elucidate the alignment brittleness phenomenon. We demonstrate that the loss landscape of large language models resembles a basin, within which models perform nearly identically and outside of which they lose all capabilities. This property enables us to derive a theoretical lower bound on performance degradation during any fine-tuning and jailbreaking attacks within certain norm constraints. We also show that the basin can be readily expanded due to the over-parameterization property of neural networks. Despite these contributions, our exploration remains preliminary. We hope our work sheds light on the alignment brittleness phenomenon and motivates large-scale studies that apply the GO optimizer during the pre-training phase of larger models at cutting-edge scales, verifying the practical utility of these theoretical guarantees and the relationship between basin size and performance degradation in general fine-tuning. See Appendix˜F for more detailed discussion.
@@ Appendix A Notations
$J_{\mathcal{D}}$ Benchmark that maps from $\mathbb{R}^{d}$ to $\mathbb{R}$.
@@ Appendix B Related Work
(a) Basic
(b) Safety
(c) Math
(d) Coding
Figure 6: The 3D version of the most-case loss landscape, using Qwen2.5-7B model.
Appendix C Why Do the Loss Landscapes of LLMs Resemble Basins?
@@
This paper presents a seemingly different conclusion from Peng et al. (2024), where we argue that the loss landscape of large language models resembles a basin, whereas they suggest that LLM performance degrades gradually as the perturbation budget increases. The key reason is that Peng et al. (2024) uses benchmarks that evaluate models by log-likelihood, where the loss landscape is expected to be continuous and smooth (See Fig.˜8(a)). In contrast, we use benchmarks that evaluate models based on their generation results.
This raises the question of why large language models perform nearly identically within a certain range of Gaussian perturbations to their parameters but rapidly lose all capabilities when perturbations exceed this range. This phenomenon aligns with recent findings that large language models can resist common noise perturbations (Men et al., 2024; Gromov et al., 2024; Sun et al., 2023), which show that removing certain modules in LLMs leaves their capabilities largely intact, but removing more causes a swift loss of conversational ability.
@@
In this work, we do not have a definitive answer to this question. However, we hypothesize that this may stem from the over-parameterization property of LLMs. Mode connectivity is an important property of over-parameterized models, stating that the modes of an over-parameterized network are connected to one another (Garipov et al., 2018; Frankle et al., 2020; Lubana et al., 2023). For large language models, beyond permutation invariance in feed-forward networks (Ainsworth et al., 2022), there is also rotational invariance in key and query matrices. Consequently, these equivalent modes may be closer to each other, forming a subspace with dimensions nearly equivalent to the original space. We leave this exploration to future work.
@@ D.2 Raw Numerical Values of Basins
Table 1: The raw benchmark value $J_{f,\mathcal{D}}(\bm{\theta}+\alpha\bm{\delta})$ ($\uparrow$, the higher the better) of the most-case landscape using Qwen2.5-7B. Due to page limitations, we present only half of the data here. Thus, the basin size is twice as large as shown, accounting for (nearly) symmetric counterpart on the other side. As shown, the loss landscape literally forms basins, within these basins, the benchmark values remain literally unchanged.
@@ D.3 Loss Landscapes of Large Models
We also visualize the loss landscapes of larger models, as shown in Fig.˜7. We draw the following conclusions:
@@
0: Benchmark values $\{J_{f,\mathcal{D}}(\bm{\theta}+\alpha\bm{\delta})\}_{\alpha\in A}$
1: Compute minimum value: $\text{min\_val}\leftarrow\min_{\alpha}J_{f,\mathcal{D}}(\bm{\theta}+\alpha\bm{\delta})$
2: Compute maximum value: $\text{max\_val}\leftarrow\max_{\alpha}J_{f,\mathcal{D}}(\bm{\theta}+\alpha\bm{\delta})$
@@
5:for each $\text{value}\in\{J_{f,\mathcal{D}}(\bm{\theta}+\alpha\bm{\delta})\}_{\alpha\in A}$ do
@@ D.4 Hypothesis Testing with Clopper-Pearson Bound
Soft Definition of Basins. To verify the soft definition of basins (Definition 3), we aim to confirm whether $J_{f,\mathcal{D}}(\bm{\theta})-\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[J_{f,\mathcal{D}}(\bm{\theta}+\bm{\epsilon})]\leq\epsilon$. We achieve this using the Clopper-Pearson bound by computing a confidence interval $p_{\text{lower}},p_{\text{upper}}]$ for $\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[J_{f,\mathcal{D}}(\bm{\theta}+\bm{\epsilon})]$ with a specified type-I error $\gamma$. Specifically, the Clopper-Pearson bound ensures that $\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[J_{f,\mathcal{D}}(\bm{\theta}+\bm{\epsilon})]\in[p_{\text{lower}},p_{\text{upper}}]$ with probability at least $1-\gamma$. If $J_{f,\mathcal{D}}(\bm{\theta})-p_{\text{lower}}\leq\epsilon$, we can assert, with type-I error $\gamma$, that the soft basin condition holds.
Strict Definition of Basins. For the strict definition of basins (Sec.˜3.2), our goal is to estimate the proportion of directions where the normalized benchmark value is exactly 0, i.e., to compute a lower bound for $\mathbb{E}_{\bm{\delta}\sim\mathcal{N}(\bm{0},\bm{I})}[\mathbb{I}\{\mathcal{T}\circ J_{f,\mathcal{D}}(\bm{\theta}+\alpha\bm{\delta})=0\}]$. The Clopper-Pearson bound provides a confidence interval $p_{\text{lower}},p_{\text{upper}}]$ with type-I error $\gamma$, allowing us to assert that at least $p_{\text{lower}}\times 100\%$ of directions form a strict basin with probability $1-\gamma$.
@@
Formulation and Application. Let $X$ denote a binomial random variable representing the number of successes in $n$ independent Bernoulli trials, where a success indicates that the perturbed model $f_{\bm{\theta}+\bm{\epsilon}}$ retains performance comparable to the original model $f_{\bm{\theta}}$. For the strict basin (Sec.˜3.2), a success occurs when $\mathcal{T}\circ J_{f,\mathcal{D}}(\bm{\theta}+\alpha\bm{\delta})=0$. For the soft basin (Eq.˜3), since $J_{f,\mathcal{D}}(\bm{\theta}+\bm{\epsilon})$ is also an expectation of the Bernoulli variable (i.e., 0-1 loss on each instance) over $\mathcal{D}$, a success occurs when the answer of a sampled instance is judged as "correct" or "safe". Given a confidence level $1-\gamma$ (we use $\gamma=0.01$), the Clopper-Pearson confidence interval $p_{\text{lower}},p_{\text{upper}}]$ for $p$ is defined by:
$P(X\geq x\mid p=p_{\text{lower}})$ $=\frac{\gamma}{2},$ (9)
$P(X\leq x\mid p=p_{\text{upper}})$ $=\frac{\gamma}{2},$ (10)
@@
$p_{\text{lower}}$ $=I^{-1}_{x,n-x+1}\left(\frac{\gamma}{2}\right),$ (11)
$p_{\text{upper}}$ $=I^{-1}_{x+1,n-x}\left(1-\frac{\gamma}{2}\right),$ (12)
@@
Experimental Settings. Following Cohen et al. (2019); Salman et al. (2019); Chen et al. (2025a), we use Monte Carlo sampling with the Clopper-Pearson bound to estimate basin sizes. We set a type-I error of $\gamma=0.01$ and a sample size of $n=100,000$ to ensure precise confidence intervals. For the soft basin, we evaluate $J_{f,\mathcal{D}}(\bm{\theta}+\bm{\epsilon})$ with $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})$ and $\sigma=0.01$, and obtain the statistical lower bound for $\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[J_{f,\mathcal{D}}(\bm{\theta}+\bm{\epsilon})]$. For the strict basin, we sample directions $\bm{\delta}\sim\mathcal{N}(\bm{0},\bm{I})$ and compute $\mathcal{T}\circ J_{f,\mathcal{D}}(\bm{\theta}+\alpha\bm{\delta})$ with $\alpha\in[0,3\times 10^{-3}]$, as in Table˜1. These settings enable reliable validation of the most-case basin properties, as visualized in Fig.˜1.
@@
Figure 8: Loss landscape and hypothesis testing results using a likelihood-based benchmark. (a) The loss landscape of Qwen2.5-7B for a likelihood-based benchmark is smooth and continuous. (b) The lower bound $p_{\text{lower}}$ from the Clopper-Pearson bound for the soft basin definition (i.e., the lower bound for $\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[J_{f,\mathcal{D}}(\bm{\theta}+\bm{\epsilon})]$) and the strict basin definition (i.e., the lower bound for proportion of directions achieving the exact zero loss).
Experimental Results. The statistical lower bounds obtained for both the strict and soft definitions of basins are shown in Fig.˜8(b). As demonstrated, the lower bound for the soft basin definition significantly exceeds that of the strict basin definition, aligning with our observation that when $\epsilon\to 0$ in Definition 3, the soft definition becomes the strict definition of basins, assuming the expected loss does not decrease after adding noise. We present the respective conclusions for the two definitions as follows:
Results for Strict Definition of Basins. In the strict definition of basins, the lower bound $p_{\text{lower}}$ represents the proportion of directions forming strict basins. As shown in Fig.˜8(b), for $\sigma=0.01$, we can assert that more than 90% of directions form strict basins, whereas for $\sigma=0.02$, we can assert that more than 65% of directions form strict basins. As $\sigma$ increases, the proportion of directions that form strict basins decreases.
Results for Soft Definition of Basins. In the soft definition of basins, the lower bound $p_{\text{lower}}$ provides a statistical guarantee for $\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[J_{f,\mathcal{D}}(\bm{\theta}+\bm{\epsilon})]$. As demonstrated, larger $\sigma$ values lead to worse model performance under noise perturbations but reduce the Lipschitz constant of the smoothed model. This reflects a classical robustness-accuracy tradeoff in randomized smoothing (Cohen et al., 2019; Salman et al., 2019; Chen et al., 2024). Choosing a larger $\sigma$ enhances model robustness but may compromise accuracy, potentially resulting in either improved or degraded performance. In Sec.˜4, we select $\sigma=0.03$, for which $p_{\text{lower}}$ remains above 0.9.

### v3

@@ 6 Conclusion
In this work, we explore the loss landscape of large language models to elucidate the alignment brittleness phenomenon. We demonstrate that the loss landscape of large language models resembles a basin, within which models perform nearly identically and outside of which they lose all capabilities. This property enables us to derive a theoretical lower bound on performance degradation during any fine-tuning and jailbreaking attacks within certain norm constraints. We also show that the basin can be readily expanded, and suppressing degradation under Gaussian noise directly translates to minimized performance degradation during subsequent fine-tuning. Despite these contributions, our exploration remains preliminary. We hope our work sheds light on the alignment brittleness phenomenon and motivates large-scale studies at cutting-edge scales. See Appendix˜F for more detailed discussion on the scalability of our method.
Acknowledgments
This work was supported by NSFC Projects (Nos. 625B2104, 62276149, 92370124, 92270001, 62350080, 92248303, U2341228, 62061136001, 62076147), the Fundamental and Interdisciplinary Disciplines Breakthrough Plan of the Ministry of Education of China (No. JYB2025XDXM101), BNRist (BNR2022RC01006), Beijing Natural Science Foundation (QY24035), CCF-BaiChuan-Ebtech Foundation Model Fund, Tsinghua Institute for Guo Qiang, and the High Performance Computing Center, Tsinghua University. J. Zhu was supported by the XPlorer Prize.
We would also like to thank Kaiyue Wen from Stanford University, Haodong Wen and Huaqing Zhang from Tsinghua University for their insightful feedback.
@@ Appendix A Notations
$\mathcal{S}_{\mathcal{D}}$ Benchmark that maps from $\mathbb{R}^{d}$ to $\mathbb{R}$.
$L$ Loss function that the smaller the better.
@@ B.1 The Resilience to Guassian Noise
B.2 On Randomized Smoothing
Randomized Smoothing (RS) was originally proposed as a statistically certified defense against adversarial examples. Cohen et al. (2019) first derived the tight certification radius for Gaussian noise using the Neyman-Pearson lemma, establishing a probabilistic guarantee for classification consistency. Subsequently, Salman et al. (2019) interpreted RS through the lens of Lipschitz continuity, demonstrating that the smoothed classifier effectively bounds the gradient norm of the function, a perspective we adopt in Eq.˜4. More recently, Chen et al. (2025a) generalized RS from Gaussian distributions to arbitrary distributions by formulating the certification as a knapsack problem, significantly expanding the applicability of the technique.
In this work, we reinterpret RS within a broader scope. The reliability of a deep learning model can be conceptually modeled as a function of three primary variables: the model parameters $\bm{\theta}$, the training data $\mathcal{D}$, and the test input $\bm{x}$. Consequently, RS serves as a certification tool by convolving one of these variables with noise. Its implication depends on the target variable:
  • RS on Input Space ($\bm{x}$): This is the standard setting (Cohen et al., 2019; Salman et al., 2019), where noise is added to the test input. This certifies robustness against Adversarial Examples, ensuring prediction stability within an $\ell_{2}$ ball around $\bm{x}$.
  • RS on Data Space ($\mathcal{D}$): Other works (e.g., Rosenfeld et al. (2020)) apply smoothing to the training data. This certifies robustness against Data Poisoning, ensuring the training outcome remains stable despite modifications to the dataset.
  • RS on Parameter Space ($\bm{\theta}$): This represents the focus of our work. We apply RS directly to the parameter space to certify robustness against Fine-tuning Degradation. This guarantees that if the parameters update within a certain radius (e.g., during benign fine-tuning), the model’s capabilities will not collapse.
In this work, we extend the application of RS to the parameter dimension. We note that the validity of this application relies on the specific geometry of the 0-1 capability landscape: it is the existence of stable basins—where discrete task performance remains invariant under noise—that ensures the smoothed model retains high performance. This stability is the prerequisite that renders parameter-space smoothing theoretically meaningful and yields non-vacuous analysis.
B.3 Relation to "The Uncanny Valley" and Confidence-Induced Flatness
Recent work by Walter et al. (2025) provides a nuanced analysis of the loss landscape, identifying the "Uncanny Valley" phenomenon where adversarial examples reside in flat but confident regions. They argue that flatness implies local but not global robustness.
We thoroughly agree with this distinction. For a classification task, global flatness is not likely possible without sacrificing utility (e.g., distinguishing classes requires non-zero gradients at decision boundaries) (Tsuzuku et al., 2018). Therefore, flatness can only theoretically guarantee robustness within a local radius.
However, we emphasize a fundamental distinction between Adversarial Attacks and Fine-tuning:
  • Adversarial Attacks seek to traverse the landscape globally to find a failure mode (potentially jumping into the "Uncanny Valley" of high-confidence errors).
  • Fine-tuning (Parameter Space) operates in a fundamentally different regime. As we discuss in Secs.˜E.5 and F.1, fine-tuning large models resembles the "Lazy Training" regime of NTK, where the parameter displacement $\|\Delta\bm{\theta}\|$ is minimal.
In this specific context of fine-tuning, the optimization trajectory is inherently constrained to a local neighborhood (Li & Zhang, 2021). As model scale increases, the "safety basin" expands (as shown in Sec.˜3.2) while the required fine-tuning displacement shrinks (see Appendix˜F). Consequently, for fine-tuning, local robustness effectively functions as global robustness, as the parameters naturally tend to remain within the basin of the pre-trained solution.
Furthermore, we concur with the derivation in Walter et al. (2025) that flatness is intrinsically linked to model confidence. Our Randomized Smoothing framework (Eq.˜5) arrives at a congruent conclusion: as the clean accuracy $p_{A}\to 1$ (high confidence), the Lipschitz constant of the smoothed classifier vanishes, inducing flatness. While Walter et al. (2025) highlight the risks of this property when predictions are wrong, our work leverages it for preservation: high confidence in correct pre-trained knowledge induces a "safety basin" that we aim to maintain. Thus, our work serves as a complementary perspective, focusing on the retention of correct capabilities rather than the genesis of adversarial errors.
Appendix C Discussions
C.1 Why Do the Loss Landscapes of LLMs Resemble Basins?
@@
This paper presents a seemingly different conclusion from Peng et al. (2024), where we argue that the loss landscape of large language models resembles a basin, whereas they suggest that LLM performance degrades gradually as the perturbation budget increases. The primary reason is the choice of metric: Peng et al. (2024) uses log-likelihood (NLL) benchmarks, where the landscape is expected to be continuous and smooth (see Fig.˜9(a)). In contrast, we use generative benchmarks that evaluate models based on discrete task success (0-1 loss), which reveals the "basin" structure hidden in the smooth likelihood surface.
To further elucidate the formation and nature of these basins, we discuss three key factors:
Semantic Stability beyond Simple Thresholding. While the discrete nature of token generation (argmax) contributes to the basin shape by creating a thresholding effect, we argue that the basin represents a deeper semantic stability. As detailed in Sec.˜D.9, we observe that within the basin, the model’s generated sentences often change structurally (e.g., different phrasings) while the final answer remains correct. This suggests the basin captures a manifold of semantically equivalent functions rather than merely a rigid region where probability rankings remain identical.
Interplay between Task Difficulty and Optimization. We observe that basin width correlates with task difficulty: simpler tasks (e.g., MMLU) tend to exhibit wider basins than harder tasks (e.g., GPQA). This phenomenon aligns with theoretical frameworks suggesting that Stochastic Gradient Descent (SGD) introduces an implicit bias towards flatter global minimizers (Damian et al., 2021; Li et al., 2021; 2025b). Simpler tasks may allow the optimization process to reach these "zero-loss" flat regions more easily or earlier during pre-training, resulting in wider, more robust basins compared to complex reasoning tasks where the solution space is sharper.
@@
Over-parameterization and Mode Connectivity. The existence of such basins may also (partially) stem from the over-parameterization of LLMs. Mode connectivity theory posits that the optima of over-parameterized networks are connected via low-loss paths (Garipov et al., 2018; Frankle et al., 2020; Lubana et al., 2023). For Transformer-based LLMs, beyond permutation invariance in feed-forward networks (Ainsworth et al., 2022), there is also rotational invariance in key and query matrices. Consequently, these equivalent modes may be closer to each other, forming a subspace with dimensions nearly equivalent to the original space. We leave this exploration to future work.
C.2 Relation to Optimization Dynamics and Saddle Points
Theoretically, since pre-trained models are likely $\epsilon$-stationary points rather than global minima, descent directions must exist. This raises a critical question: Despite the theoretical existence of descent directions, why are they seemingly absent in the landscape visualization, and does their existence imply a flaw in our basin-based theoretical framework? We address this by clarifying the distinction between optimization dynamics and capability geometry, and by demonstrating that our guarantees hold regardless of local optimality.
Descent Directions and Local Minima.
First, we strictly acknowledge that the pre-trained model $\theta_{0}$ has not reached a global or strict local minimum. Empirical evidence shows that SGD optimization typically converges to an $\epsilon$-stationary point where the gradient norm diminishes but does not vanish. Furthermore, theoretical analysis of the Hessian spectrum in LLMs suggests the presence of negative eigenvalues (Zhang et al., 2024c), which mathematically confirms that specific directions for further loss reduction exist. The capability of subsequent fine-tuning to further improve performance serves as practical proof that the weights preserve the potential for optimization.
Saddle Points in 0-1 vs. NLL Landscapes.
However, the concept of saddle points is typically discussed in the context of optimization dynamics on the continuous negative log-likelihood (NLL) landscape. In contrast, our work focuses on the geometry of the converged capability landscape (i.e., task success/failure). As shown in Fig.˜9, the discrete nature of 0-1 capability benchmarks often forms flat plateaus—or basins—even when the underlying NLL landscape remains curved. In this discrete setting, the optimization-centric view of saddle points does not directly contradict the observed stability of model capabilities.
Sparsity and Robustness Upper Bounds.
The apparent absence of these descent directions in our visualization is explained by the sparsity of high-dimensional space. Our method uses Gaussian noise to statistically represent "most-case" directions. In the massive parameter space of LLMs, the specific descent directions (corresponding to gradients or negative curvature) are extremely sparse; a random Gaussian vector is highly likely to be orthogonal to them, resulting in the observed stability. Crucially, this does not invalidate our theory, as our framework relies on robustness rather than optimality. Random Gaussian noise acts as an empirical upper bound for the degradation caused by benign fine-tuning. Since benign SFT aims to enhance the model by following those specific descent directions, it should theoretically be less harmful to existing capabilities than blind random noise. Consequently, our theoretical bound remains valid for any fine-tuning process, including those that successfully reduce the loss along specific descent directions.
@@ D.2 Raw Numerical Values of Basins
Table 1: The raw benchmark value $\mathcal{S}_{f,\mathcal{D}}(\bm{\theta}+\alpha\bm{\delta})$ ($\uparrow$, the higher the better) of the most-case landscape using Qwen2.5-7B. Due to page limitations, we present only half of the data here. Thus, the basin size is twice as large as shown, accounting for (nearly) symmetric counterpart on the other side. As shown, the loss landscape literally forms basins, within these basins, the benchmark values remain literally unchanged.
@@ D.3 Loss Landscapes of Large Models
We also visualize the loss landscapes of larger models, as shown in Fig.˜8. We draw the following conclusions:
@@
0: Benchmark values $\{\mathcal{S}_{f,\mathcal{D}}(\bm{\theta}+\alpha\bm{\delta})\}_{\alpha\in A}$
1: Compute minimum value: $\text{min\_val}\leftarrow\min_{\alpha}\mathcal{S}_{f,\mathcal{D}}(\bm{\theta}+\alpha\bm{\delta})$
2: Compute maximum value: $\text{max\_val}\leftarrow\max_{\alpha}\mathcal{S}_{f,\mathcal{D}}(\bm{\theta}+\alpha\bm{\delta})$
@@
5:for each $\text{value}\in\{\mathcal{S}_{f,\mathcal{D}}(\bm{\theta}+\alpha\bm{\delta})\}_{\alpha\in A}$ do
@@ D.4 Hypothesis Testing with Clopper-Pearson Bound
Soft Definition of Basins. To verify the soft definition of basins (Definition 3), we aim to confirm whether $\mathcal{S}_{f,\mathcal{D}}(\bm{\theta})-\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[\mathcal{S}_{f,\mathcal{D}}(\bm{\theta}+\bm{\epsilon})]\leq\tau$. We achieve this using the Clopper-Pearson bound by computing a confidence interval $p_{\text{lower}},p_{\text{upper}}]$ for $\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[\mathcal{S}_{f,\mathcal{D}}(\bm{\theta}+\bm{\epsilon})]$ with a specified type-I error $\gamma$. Specifically, the Clopper-Pearson bound ensures that $\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[\mathcal{S}_{f,\mathcal{D}}(\bm{\theta}+\bm{\epsilon})]\in[p_{\text{lower}},p_{\text{upper}}]$ with probability at least $1-\gamma$. If $\mathcal{S}_{f,\mathcal{D}}(\bm{\theta})-p_{\text{lower}}\leq\epsilon$, we can assert, with type-I error $\gamma$, that the soft basin condition holds.
Strict Definition of Basins. For the strict definition of basins (Sec.˜3.2), our goal is to estimate the proportion of directions where the normalized benchmark value is exactly 0, i.e., to compute a lower bound for $\mathbb{E}_{\bm{\delta}\sim\mathcal{N}(\bm{0},\bm{I})}[\mathbb{I}\{\mathcal{T}\circ\mathcal{S}_{f,\mathcal{D}}(\bm{\theta}+\alpha\bm{\delta})=0\}]$. The Clopper-Pearson bound provides a confidence interval $p_{\text{lower}},p_{\text{upper}}]$ with type-I error $\gamma$, allowing us to assert that at least $p_{\text{lower}}\times 100\%$ of directions form a strict basin with probability $1-\gamma$.
@@
Formulation and Application. Let $X$ denote a binomial random variable representing the number of successes in $n$ independent Bernoulli trials, where a success indicates that the perturbed model $f_{\bm{\theta}+\bm{\epsilon}}$ retains performance comparable to the original model $f_{\bm{\theta}}$. For the strict basin (Sec.˜3.2), a success occurs when $\mathcal{T}\circ\mathcal{S}_{f,\mathcal{D}}(\bm{\theta}+\alpha\bm{\delta})=0$. For the soft basin (Eq.˜3), since $\mathcal{S}_{f,\mathcal{D}}(\bm{\theta}+\bm{\epsilon})$ is also an expectation of the Bernoulli variable (i.e., 0-1 loss on each instance) over $\mathcal{D}$, a success occurs when the answer of a sampled instance is judged as "correct" or "safe". Given a confidence level $1-\gamma$ (we use $\gamma=0.01$), the Clopper-Pearson confidence interval $p_{\text{lower}},p_{\text{upper}}]$ for $p$ is defined by:
$P(X\geq x\mid p=p_{\text{lower}})=\frac{\gamma}{2},\;P(X\leq x\mid p=p_{\text{upper}})=\frac{\gamma}{2}$
@@
$p_{\text{lower}}=I^{-1}_{x,n-x+1}\left(\frac{\gamma}{2}\right),\;p_{\text{upper}}=I^{-1}_{x+1,n-x}\left(1-\frac{\gamma}{2}\right),$
@@
Experimental Settings. Following Cohen et al. (2019); Salman et al. (2019); Chen et al. (2025a), we use Monte Carlo sampling with the Clopper-Pearson bound to estimate basin sizes. We set a type-I error of $\gamma=0.01$ and a sample size of $n=100,000$ to ensure precise confidence intervals. For the soft basin, we evaluate $\mathcal{S}_{f,\mathcal{D}}(\bm{\theta}+\bm{\epsilon})$ with $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})$ and $\sigma=0.01$, and obtain the statistical lower bound for $\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[\mathcal{S}_{f,\mathcal{D}}(\bm{\theta}+\bm{\epsilon})]$. For the strict basin, we sample directions $\bm{\delta}\sim\mathcal{N}(\bm{0},\bm{I})$ and compute $\mathcal{T}\circ\mathcal{S}_{f,\mathcal{D}}(\bm{\theta}+\alpha\bm{\delta})$ with $\alpha\in[0,3\times 10^{-3}]$, as in Table˜1. These settings enable reliable validation of the most-case basin properties, as visualized in Fig.˜1.
@@
Figure 9: Loss landscape and hypothesis testing results using a likelihood-based benchmark. (a) The loss landscape of Qwen2.5-7B for a likelihood-based benchmark is smooth and continuous. (b) The lower bound $p_{\text{lower}}$ from the Clopper-Pearson bound for the soft basin definition (i.e., the lower bound for $\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[\mathcal{S}_{f,\mathcal{D}}(\bm{\theta}+\bm{\epsilon})]$) and the strict basin definition (i.e., the lower bound for proportion of directions achieving the exact zero loss).
Experimental Results. The statistical lower bounds obtained for both the strict and soft definitions of basins are shown in Fig.˜9(b). As demonstrated, the lower bound for the soft basin definition significantly exceeds that of the strict basin definition, aligning with our observation that when $\tau\to 0$ in Definition 3, the soft definition becomes the strict definition of basins, assuming the expected loss does not decrease after adding noise. We present the respective conclusions for the two definitions as follows:
Results for Strict Definition of Basins. In the strict definition of basins, the lower bound $p_{\text{lower}}$ represents the proportion of directions forming strict basins. As shown in Fig.˜9(b), for $\sigma=0.01$, we can assert that more than 90% of directions form strict basins, whereas for $\sigma=0.02$, we can assert that more than 65% of directions form strict basins. As $\sigma$ increases, the proportion of directions that form strict basins decreases.
Results for Soft Definition of Basins. In the soft definition of basins, the lower bound $p_{\text{lower}}$ provides a statistical guarantee for $\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[\mathcal{S}_{f,\mathcal{D}}(\bm{\theta}+\bm{\epsilon})]$. As demonstrated, larger $\sigma$ values lead to worse model performance under noise perturbations but reduce the Lipschitz constant of the smoothed model. This reflects a classical robustness-accuracy tradeoff in randomized smoothing (Cohen et al., 2019; Salman et al., 2019; Chen et al., 2024). Choosing a larger $\sigma$ enhances model robustness but may compromise accuracy, potentially resulting in either improved or degraded performance. In Sec.˜4, we select $\sigma=0.03$, for which $p_{\text{lower}}$ remains above 0.9.